Grade 10Mechanics

The sum of two numbers is 6. If one is twice the other, find the numbers.

samiullah
9 Years agoGrade 10

2 Answers

Vikas TU
9 Years ago
 
let the two numbers be x and respectively.
And given,
x = 2y
and x + y = 6
solving both the eqns. we get,
3y = 6
y = 2
and
x = 4
Rishabh singh
9 Years ago
Let one number be x, And another no be y=2x (given), So a.t.q (ACCORDING TO QUESTION),, x+y=6, (given in the question) x+2x=6, 3x=6, so, x=2, and y=4